# Adapting a novel framework for rapid inference of massive black hole binaries for LISA

### v1

@@ III.2 LISA data challenge
Fig 1 indicates that signal-1 merges one day after the merger of signal-0. To investigate the impact of nearby MBHB signals on the recovery of signal-0, we first carry out zero-noise inference of signal-0 in three scenarios: in scenario 1 the data only contains signal-0, in scenario 2 the data contains signal-0 and signal-1, and in scenario 3 the data contains all six signals. For this investigation, we generate the PSDs by applying the Welch method to the LDC-2A training dataset, excluding GBs from its generation. The PSD for $A$ channel is plotted in blue in Fig 3, and using these PSDs the SNR of signal-0 is approximately $1277$.
@@
Figure 4: Zero-noise injection recovery of signal-0 in multiple scenarios: This figure illustrates the results of analyzing signal-0 in three scenarios: scenario 1 only has signal-0 in the data (blue), scenario 2 has signal-0 and signal-1 in the data (orange), and scenario 3 has all six signals in the data (green). For both plots, diagonal panels show the one-dimensional marginal posterior distribution, while contours in the off-diagonal panels show the 90% credible intervals for the two-dimensional marginal posterior distribution. The crosshairs and vertical lines mark the true values. We find that the presence of other signals does not significantly affect the analysis of signal-0.
The results of this investigation are illustrated in Fig 4. From the figure, we can see that the introduction of other signals has minimal impact on the recovery of the $m_{1},m_{2},a_{1z}$ and $a_{2z}$. However, the impact is more pronounced if we express the masses as ${\cal M}_{c}$. To quantify the differences in the posteriors from scenarios 2 and 3 when compared to posteriors from scenario 1, we calculate absolute normalized bias. It is defined as $|\Delta\tilde{\lambda}|/\sigma$ where $\Delta\tilde{\lambda}$ is the difference in the median of $\lambda$ posterior distribution and $\sigma$ is the standard deviation of $\lambda$ posterior from scenario 1. For both scenarios 2 and 3, the normalized bias in $m_{1},m_{2},a_{1z}$ and $a_{2z}$ is much lower than $0.15$. In contrast, for ${\cal M}_{c}$, the bias is around $0.55$. While a bias greater than $1$ is typically seen as significant, this value is notably larger than those obtained for the fundamental mass and spin parameters. The bias in $D_{L}$ is approximately $0.35$ for both scenarios.
@@
Figure 5: LDC-2A (Sangria) analysis: This figure illustrates the results of our analysis of signal-0 in the LDC-2A dataset which contains six MBHB signals, instrumental noise, and GBs. For both plots, diagonal panels show the one-dimensional marginal posterior distribution, while contours in the off-diagonal panels show the 90% credible intervals for the two-dimensional marginal posterior distribution. In the top plot, the colored points represent the $\text{ln}\mathcal{L}_{\text{marg}}$ values used in constructing the posteriors, within the range of $\Delta\text{ln}\mathcal{L}_{\text{marg}}\leq 15$ measured with respect to the maximum $\text{ln}\mathcal{L}_{\text{marg}}$ value. The crosshairs and vertical lines mark the true values. We find that the presence of GBs and noise significantly impacts the inference of signal-0 by shifting the posteriors away from the true value.
@@
Radler: From the Radler challenge, we analyze the LDC-1A dataset which includes a single MBHB signal with mass and spin parameters listed in Table 3. This dataset spans nearly 1.33 years and is sampled at ten-second intervals. We analyze the MBHB signal at a zero-noise realization and use “SciRDv1” PSDs in our likelihood computations. With these PSDs, we find the SNR of the MBHB signal to be approximately $390$. Fig 6 illustrates the results of our analysis, indicating that the true values are well within the $90\%$ credible interval. We observe that our analysis constrains the mass and spin parameters more tightly than Katz (2022). Additionally, there is a slight shift in the peak for mass posteriors relative to the true values. Further investigation revealed that the ${\cal M}_{c}$ posterior peaks at the true value, but $\eta$ posterior was slightly shifted, leading to corresponding shifts in the mass posteriors. The color gradient suggests that this shift stems from the underlying marginalized likelihood values. This could potentially be caused by differences in data conditioning and/or due to the differences in the likelihood method used; Katz (2022) used heterodyned likelihood, while our analysis utilized full Whittle likelihood.
@@
Figure 6: LDC-1A (Radler) analysis: This figure illustrates the results of our analysis of the LDC-1A zero-noise dataset. Diagonal panels show the one-dimensional marginal posterior distribution, while contours in the off-diagonal panels show the 90% credible intervals for the two-dimensional marginal posterior distribution. The colored points represent the $\text{ln}\mathcal{L}_{\text{marg}}$ values used in constructing the posteriors, within the range of $\Delta\text{ln}\mathcal{L}_{\text{marg}}\leq 15$ measured with respect to the maximum $\text{ln}\mathcal{L}_{\text{marg}}$ value. The crosshairs and vertical lines mark the true values.
@@ IV Discussion
Motivated by the need for a code capable of analyzing long-duration, high SNR signals using models that might be prohibitive to use with standard sampling techniques, we modify the RIFT parameter estimation algorithm for LISA. This modification, called LISA-RIFT, takes advantage of RIFT’s embarrassingly parallel architecture, enabling the use of costly waveform models at full Fourier bin resolution. We find LISA-RIFT’s computational performance to be highly efficient in terms of both wall clock and core time. For instance, the IMRPhenomD analysis shown in Fig 4 only took around $10$ hours of wall clock time and approximately $350$ core-days ($\simeq 10^{4}$ core-hours) of CPU time. The bulk of the computational cost came from around $10^{5}$ marginalized likelihood evaluations performed during this analysis, each taking about $30$ seconds. This analysis was performed on a year-long dataset at full Fourier bin resolution. In comparison, standard sampling techniques analyzing the same dataset at full Fourier bin resolution would require approximately $O(10^{5})$ CPU hours, ten times more than our approach. Additionally, our approach could be further accelerated by leveraging RIFT’s native hardware accleration Wysocki et al. (2019). A similar performance was observed in our analysis using NRHybSur3dq8. That analysis, performed on a four month long dataset with both injection and recovery involving nine different GW modes, took around 20 hours of wall clock time.
@@ Acknowledgements.
We would like to thank Connor Weaving for his help with the LDC-2A dataset. We thank Sylvian Marsat for addressing our questions regarding the LISA response model, and Mike Katz for making his BBHx code publicly available, which was used to validate our implementation of the response model. This research was done using services provided by the OSG Consortium Pordes et al. (2007); Sfiligoi et al. (2009); OSG (2006, 2015) which is supported by the National Science Foundation awards 2030508 and 1836650. Parts of this code were tested on LLO and LHO cluster of the LIGO scientific collaboration supported by NSF Grants PHY-0757058 and PHY-0823459. AJ and DS acknowledge support from NASA 80NSSC24K0437 and NSF Grants PHY-2207780 and PHY-2114581. ROS gratefully acknowledges support from NSF awards NSF PHY-1912632, PHY-2012057, PHY-2309172, AST-2206321, and the Simons Foundation.

### v2

@@ III.2 LISA data challenge
Figure 1 indicates that signal-1 merges one day after the merger of signal-0. To investigate the impact of nearby MBHB signals on the recovery of signal-0, we first carry out zero-noise inference of signal-0 in three scenarios: in scenario 1 the data only contains signal-0, in scenario 2 the data contains signal-0 and signal-1, and in scenario 3 the data contains all six signals. For this investigation, we generate the PSDs by applying the Welch method to the LDC-2A training dataset, excluding GBs from its generation. The PSD for $A$ channel is plotted in blue in Fig 3, and using these PSDs the SNR of signal-0 is approximately $1277$.
@@
Figure 4: Zero-noise injection recovery of signal-0 in multiple scenarios: This figure illustrates the results of analyzing signal-0 in three scenarios: scenario 1 only has signal-0 in the data (blue), scenario 2 has signal-0 and signal-1 in the data (orange), and scenario 3 has all six signals in the data (green). For both plots, diagonal panels show the one-dimensional marginal posterior distribution, while contours in the off-diagonal panels show the 90% credible intervals for the two-dimensional marginal posterior distribution. The crosshairs and vertical lines mark the true values. In all three scenarios, the dataset is one year long. We find that the presence of other signals does not significantly affect the analysis of signal-0.
The results of this investigation are illustrated in Fig 4. From the figure, we can see that the introduction of other signals has minimal impact on the recovery of the $m_{1},m_{2},a_{1z}$ and $a_{2z}$. However, the impact is more pronounced if we express the masses as ${\cal M}_{c}$. To quantify the differences in the posteriors from scenarios 2 and 3 when compared to posteriors from scenario 1, we calculate absolute normalized bias. It is defined as $|\Delta\tilde{\kappa}|/\sigma$ where $\Delta\tilde{\kappa}$ is the difference in the median of $\kappa$ posterior distributions and $\sigma$ is the standard deviation of $\kappa$ posterior from scenario 1. For both scenarios 2 and 3, the normalized bias in $m_{1},m_{2},a_{1z}$ and $a_{2z}$ is much lower than $0.15$. In contrast, for ${\cal M}_{c}$, the bias is around $0.55$. While a bias greater than $1$ is typically seen as significant, this value is notably larger than those obtained for the fundamental mass and spin parameters. The bias in $D_{L}$ is approximately $0.35$ for both scenarios.
@@
Figure 5: LDC-2A (Sangria) analysis: This figure illustrates the results of our analysis of signal-0 in the year long LDC-2A dataset which contains six MBHB signals, instrumental noise, and GBs. For both plots, diagonal panels show the one-dimensional marginal posterior distribution, while contours in the off-diagonal panels show the 90% credible intervals for the two-dimensional marginal posterior distribution. In the top plot, the colored points represent the $\text{ln}\mathcal{L}_{\text{marg}}$ values used in constructing the posteriors, within the range of $\Delta\text{ln}\mathcal{L}_{\text{marg}}\leq 15$ measured with respect to the maximum $\text{ln}\mathcal{L}_{\text{marg}}$ value. The crosshairs and vertical lines mark the true values. We find that the presence of GBs and noise significantly impacts the inference of signal-0 by shifting the posteriors away from the true value.
@@
Radler: From the Radler challenge, we analyze the LDC-1A dataset which includes a single MBHB signal with mass and spin parameters listed in Table 3. This dataset spans nearly 1.33 years and is sampled at ten-second intervals. We analyze the MBHB signal at a zero-noise realization and use “SciRDv1” PSDs in our likelihood computations. With these PSDs, we find the SNR of the MBHB signal to be approximately $390$. Figure 6 illustrates the results of our analysis, indicating that the true values are well within the $90\%$ credible interval. We observe that our analysis constrains the mass and spin parameters more tightly than Katz (2022). Additionally, there is a slight shift in the peak for mass posteriors relative to the true values. Further investigation revealed that the ${\cal M}_{c}$ posterior peaks at the true value, but $\eta$ posterior was slightly shifted, leading to corresponding shifts in the mass posteriors. The color gradient suggests that this shift stems from the underlying marginalized likelihood values. This could potentially be caused by differences in data conditioning and/or due to the differences in the likelihood method used; Katz (2022) used heterodyned likelihood, while our analysis utilized full Whittle likelihood.
@@
Figure 6: LDC-1A (Radler) analysis: This figure illustrates the results of our analysis of the 1.33 years long LDC-1A zero-noise dataset. Diagonal panels show the one-dimensional marginal posterior distribution, while contours in the off-diagonal panels show the 90% credible intervals for the two-dimensional marginal posterior distribution. The colored points represent the $\text{ln}\mathcal{L}_{\text{marg}}$ values used in constructing the posteriors, within the range of $\Delta\text{ln}\mathcal{L}_{\text{marg}}\leq 15$ measured with respect to the maximum $\text{ln}\mathcal{L}_{\text{marg}}$ value. The crosshairs and vertical lines mark the true values.
@@ IV Discussion
Motivated by the need for a code capable of analyzing long-duration, high SNR signals using models that might be prohibitive to use with standard sampling techniques, we modify the RIFT parameter estimation algorithm for LISA. This modification, called LISA-RIFT, takes advantage of RIFT’s embarrassingly parallel architecture, enabling the use of costly waveform models at full Fourier bin resolution. We find LISA-RIFT’s computational performance to be highly efficient in terms of both wall clock and core time. For instance, the IMRPhenomD analysis shown in Fig 4 and Fig 5 only took around $10$ hours of wall clock time and approximately $350$ core-days ($\approx 10^{4}$ core-hours) of CPU time. The bulk of the computational cost came from around $10^{5}$ marginalized likelihood evaluations performed during this analysis, each taking about $30$ seconds. This analysis was performed on a year-long dataset at full Fourier bin resolution. In comparison, standard sampling techniques using heterodyned likelihood when applied to LDC datasets typically take anywhere from a few hours Cornish and Shuman (2020) to nearly half a day Weaving et al. (2023) of wall clock time to complete their analysis. Considering the $10^{2}$ speedup provided by heterodyning, these methods would require anywhere from a few days to a month to complete when utilizing the full Whittle likelihood. Additionally, our approach could be further accelerated by leveraging RIFT’s native hardware acceleration Wysocki et al. (2019). A similar performance was observed in our analysis using NRHybSur3dq8. That analysis, performed on a four month long dataset with both injection and recovery involving nine different GW modes, took around 20 hours of wall clock time.
@@ Acknowledgements.
We would like to thank Connor Weaving for his help with the LDC-2A dataset. We thank Sylvian Marsat for addressing our questions regarding the LISA response model, and Michael Katz for making his BBHx code publicly available, which was used to validate our implementation of the response model. We thank the anonymous reviewer for their helpful comments. This research was done using services provided by the OSG Consortium Pordes et al. (2007); Sfiligoi et al. (2009); OSG (2006, 2015) which is supported by the National Science Foundation awards No. 2030508 and No. 1836650. Parts of this code were tested on LLO and LHO cluster of the LIGO scientific collaboration supported by NSF Grants No. PHY-0757058 and No. PHY-0823459. A.J. and D.S. acknowledge support from NASA Grant No. 80NSSC24K0437. A.J., D.S., and J.L. acknowledge support from NSF Grants No. PHY-2207780 and No. PHY-2114581. R.O.S. gratefully acknowledges support from NSF awards No. NSF PHY-1912632, No. PHY-2012057, No. PHY-2309172, No. AST-2206321, and the Simons Foundation. The work was done by members of the Weinberg Institute and has an identifier of UT-WI-36-2024.